! Required information The Foundational 15 (Algo) (LO5-1, LO5-3, LO5-4, LO5-5, LO5-6, LO5-7, LO5-8] [The following informatio11. What is the margin of safety in dollars? What is the margin of safety percentage? Margin of safety in dollars Margin of s12. What is the degree of operating leverage? (Round your answer to 2 decimal places.) Degree of operating leverage13. Using the degree of operating leverage, what is the estimated percent increase in net operating income of a 5% increase i14. Assume that the amounts of the companys total variable expenses and total fixed expenses were reversed. In other words,15. Assume that the amounts of the companys total variable expenses and total fixed expenses were reversed. In other words,
Answer #1

11.

Contribution Margin ratio = Contribution margin / sales = $22,500 / $50,000 = 0.45

Breakeven point (sales in dollars) = Fixed cost / Contribution margin ratio = $14,850 / 0.45 = $33,000

Margin of safety = Actual sales - Break even sales = $50,000 - $33,000 = $ 17,000

Margin of safety percentage = (Actual sales - Break even sales) / Actual sales

= $17,000 / $50,000 = 0.34

Margin of safety percentage = 34%

12.

Degree of operating leverage = Contribution margin/net operating income

= 22,500/7650 = 2.94

13.

% increase in net operating income = Degree of operating leverage X % increase in sales

                                                                      =2.9411*5

                                                                      = 14.71%

14.

Contribution margin = $50,000 - $14,850 = $35150

Net operating income = $35,150 - $27,500 = $7650

Degree of operating leverage = $35,150 / $7650 = 4.59

15.

% increase in net operating income = Degree of operating leverage X % increase in sales

                                                                      =4.5947*5

                                                                      = 22.97 %